Ayden’s Toys, Inc., just purchased a $470,000 machine to produce toy cars. The machine will be fully depreciated by the straight-line method over its 6-year economic life. Each toy sells for $18. The variable cost per toy is $6 and the firm incurs fixed costs of $330,000 per year. The corporate tax rate for the company is 22 percent. The appropriate discount rate is 10 percent. What is the financial break-even point for the project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Depreciation per year =470000/6 =78333.33

Payment per year =PMT(10%,6,-470000) =107915.47

financial break-even =(330000+(107915.47-(78333.33*0.22))/(1-0.22))/(18-6)